loverofcats:

Select all that apply. Which steps are involved in multiplying fractions? Find a common denominator Multiply the numerators together Check to see if the product can be simplified Multiply the denominators together Convert the product to a mixed number

loverofcats:

please help and hurry

yeyedontplay:

2/3 * 2/4 = 4/12 = 1/3 Firstly you multiply the Numerators. Secondly you multiply the denominators. Check to see if the product can be simplified Convert the product to a mixed number if necessary